# Refinement of Monocular Depth Maps via Multi-View Differentiable Rendering

### v1

@@ 4 Depth Refinement with GPU Acceleration
We convert the previously used intrinsics definition $C$ with $\{f_{x},f_{y},c_{x},c_{y}\}$ to the projection matrix $P$ following the OpenGL convention [2] and using homogeneous 4D coordinates for all vertices. This $P_{\{0..N\}}$ is then used in place of $C_{\{0..N\}}$ in Equ. 3.
@@
$\small z^{\prime}=\frac{d_{\textit{far}}+d_{\textit{near}}}{d_{\textit{near}}-d_{\textit{far}}}+\frac{2d_{\textit{far}}d_{\textit{near}}}{-z^{g}(d_{\textit{near}}-d_{\textit{far}})}.$ (14)
$u$ and $v$ are stored in normalized image coordinates in the $[-1,1]$ range with
@@
An upside of this normalization is that it improves numerical stability and is also simplifies positional MLP encoding (see Sec. 3.6).
@@
For both global and local refinement, depth optimization in reciprocal space leads to instable optimization. As such we do this optimization in linear space by evaluating and substitute the rightmost part in Equ. 3 first by
@@
In this representation, $\Delta u$ and $\Delta v$ are now limited by $\frac{j}{W}$ and $\frac{j}{H}$ respectively.
@@
$\small\begin{array}[]{c}p^{\prime}=\min(-\frac{d_{\textit{far}}-\epsilon}{z},1)\max(-\frac{d_{\textit{near}}+\epsilon}{z},1)p^{\prime\prime}.\end{array}$ (17)
@@
Thus, Equ. 3 can be rewritten as:
$\small p_{i}=P_{i}(R_{i}|t_{i})(R_{0}|t_{0})^{-1}p^{\prime},$ (18)
with $(R|t)$ matrices now being $4\times 4$ dimensional, respectively.
5 Results
We evaluate our approach on the synthetic Replica dataset [46, 38] and the real-world ScanNet++ dataset [67]. Both data sets provide metric scale ground truths.
For the Replica dataset, color and ground truth depth maps are obtained from the rendered meshes. We selected four scenes, and optimized 64 depth maps for each. As sparse point clouds, we generate point samples from the ground truth meshes and add 2-10% noise as well as outliers to them (see Appendix B for details).
As for the ScanNet++ dataset, ground truth depth maps are rendered from the meshes, the sparse point cloud and poses have been computed by COLMAP [40]. For the results present in Sec. 5.1, we used 10 scenes from the ScanNet++ validation set and 10 depth maps of each (100 in total). Details on the dataset preparations can be found in Appendix B. For hyperparameter tuning, we used 7 different scenes.
Implementation details, the following results were created with, are in the supplements.
5.1 Comparison against Prior Arts
We compare our approach with the top three methods of the Tanks and Temples benchmark at the time [25]. All of them are MVSNet-based methods [65]: MVSFormer [3], MVSFormer++ [5], and GeoMVSNet [74]. For GeoMVSNet however, only weights trained on the DTU dataset were available, which impacted generalization capabilities. These methods produce a certainty value that can be thresholded to control outlier removal and depth map completeness. For comparisons, we considered only depth values with valid ground truth within a 7-meter range, to stay within the specified bounding volumes for related methods, and masked values according to their certainty thresholds of 0, 0.25, and 0.5. The metrics of resulting depth maps, which vary in completeness, are documented in the tables.
As an established representative of patch match-based methods, we also compared with photometric depth maps generated by COLMAP [41]. Finally, we compared with the recent indoor-specialized metric configuration of the monocular depth estimator DepthAnything [63]. Our results were produced using hyperparameters as described in 5.2 and the supplemental. For initial depth estimations we use Marigold [21].
We consider the standard error metrics root mean squared error (RMSE), mean absolute error (MAE), L1-rel [50], L1-inv [50], $\delta_{1}$ [63]. We also calculate the accuracy of estimated depth samples ($|z-z_{GT}|<\tau$) considering different thresholds $\tau$ Results with values in the upper 20 percent are marked in green, the greener the better.
Method Valid RMSE MAE L1-rel L1-inv $\delta_{1}$ Acc. Acc. Acc.
Samples 0.01 m 0.05 m 0.10 m
MVSFormer (0) 1.00 0.80 0.31 0.13 0.17 0.88 0.21 0.58 0.73
MVSFormer++ (0) 1.00 0.56 0.27 0.10 0.22 0.89 0.23 0.61 0.76
MVSFormer (0.25) 0.97 0.75 0.27 0.11 0.11 0.90 0.21 0.58 0.73
MVSFormer++ (0.25) 0.95 0.42 0.17 0.06 0.08 0.93 0.23 0.61 0.76
MVSFormer (0.5) 0.87 0.46 0.14 0.06 0.03 0.96 0.21 0.57 0.71
MVSFormer++ (0.5) 0.85 0.19 0.07 0.03 0.01 0.98 0.22 0.59 0.73
Ours (+) 0.99 0.08 0.04 0.02 0.01 1.00 0.37 0.81 0.92
Ours (-) 0.99 0.09 0.05 0.02 0.01 0.99 0.26 0.69 0.87
Table 1: Results on the Replica dataset. In brackets are the confidence thresholding for MVSNet-based methods and sparse point cloud accuracy with good (+) and poor (-) quality.
Overall, our approach performed best for the synthetic Replica dataset, as shown in Tab. 1, independently of the quality of the point cloud. Ours+ refers to good point clouds (2 % noise + 2 % outliers), Ours- to poor point clouds (5-10 % noise + 5-10 % outliers, see supplemental material for further details). While surprisingly robust, differences to ground truth are notable especially with very poor point cloud input.
Method Valid RMSE MAE L1-rel L1-inv $\delta_{1}$ Acc. Acc. Acc.
Samples 0.01 m 0.05 m 0.10 m
COLMAP 69% 0.60 0.30 0.24 0.49 0.73 0.17 0.35 0.41
DepthAnything 100% 1.91 1.79 1.51 0.62 0.00 0.00 0.00 0.00
MVSFormer (0) 100% 1.08 0.47 0.41 0.50 0.68 0.16 0.40 0.52
MVSFormer++ (0) 100% 0.67 0.28 0.24 0.33 0.79 0.22 0.53 0.65
GeoMVSNet (0) 100% 1.19 0.61 0.77 0.36 0.53 0.05 0.18 0.28
MVSFormer (0.25) 90% 1.09 0.46 0.38 0.38 0.72 0.16 0.40 0.51
MVSFormer++ (0.25) 88% 0.61 0.22 0.19 0.14 0.86 0.22 0.52 0.63
GeoMVSNet (0.25) 100% 1.19 0.61 0.77 0.36 0.53 0.05 0.18 0.28
MVSFormer (0.5) 47% 0.52 0.16 0.10 0.06 0.93 0.14 0.30 0.36
MVSFormer++ (0.5) 62% 0.26 0.10 0.07 0.04 0.96 0.19 0.43 0.50
GeoMVSNet (0.5) 23% 1.45 0.90 1.11 0.48 0.38 0.00 0.02 0.04
Ours 100% 0.16 0.09 0.07 0.10 0.93 0.21 0.58 0.76
Table 2: Results on the Scannet++ dataset [67]. Numbers in brackets indicate the confidence thresholding for MVSNet methods.
Quantitative results for Scannet++ are shown in Tab. 2. Regarding completeness (valid samples), RMSE, MAS, L1-rel, and Accuracy $\tau$=0.05 and $\tau$=0.1, we performed best or on par overall. Only for L1-inv, $\delta_{1}$, and accuracy with $\tau$ = 0.01, MVSFormer++ with certainty threshold of 0.5 performed slightly better. However, for this configuration only 62% of the values were actually considered for the mean metric, while our results were evaluated on nearly all available samples ($>$ 99.5 %).
Figure 5: Qualitative results on the Replica dataset. Certainty threshold of MVSFormers was set to the their default of 0.5.
Qualitative results of both the Replica and the ScanNet++ dataset are shown in Figs. 5 and 6. Our results are clean, dense and plausible. However, artifact are notable at the glossy table in Scene 3. Sometimes our solutions are even more complete than the ground truth, for examples see ventilator and lamp (Scene 4, left and right crop) or, the leg of the table (Scene5, right crop). Results of MVSNet-likes are sparse and especially walls appear noisy. The metric version of DepthAnything shows very promising geometric details, however, the global scale is inaccurate.
Figure 6: Qualitative results on the ScanNet++ dataset. Certainty threshold of MVSFormers was set to the their default of 0.5.
5.2 Ablation Studies
In this section, we finetune hyper parameters and present ablation studies based on seven scenes of the Scannet++ dataset. Further ablation studies are in the Appendix C.
Losses.
Setup RMSE MAE L1-rel L1-inv Acc. Acc. Acc.
0.01 m 0.05 m 0.10 m
$\scriptstyle\mathcal{L}_{\textit{geo}}$ 0.32 0.13 0.15 0.15 0.14 0.45 0.65
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{L}_{\textit{photo}}$ 0.86 0.37 0.42 0.27 0.17 0.39 0.51
$\scriptstyle\mathcal{L}_{\textit{geo}}+{\mathcal{R}_{\textit{s}}}$ 0.26 0.12 0.12 0.15 0.14 0.45 0.65
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{R}_{\textit{p}}$ 0.22 0.11 0.12 0.12 0.16 0.53 0.74
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{R}_{\textit{p}}+\mathcal{R}_{\textit{e}}$ 0.21 0.10 0.11 0.10 0.16 0.55 0.76
$\scriptstyle\mathcal{L}_{\textit{geo}}+{\mathcal{R}_{\textit{s}}}+\mathcal{R}_{\textit{p}}+\mathcal{R}_{\textit{e}}$ 0.21 0.10 0.11 0.10 0.16 0.55 0.76
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{L}_{\textit{photo}}+\mathcal{R}_{\textit{p}}+\mathcal{R}_{\textit{e}}$ 0.22 0.10 0.12 0.13 0.22 0.57 0.75
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{L}_{\textit{photo}}+{\mathcal{R}_{\textit{s}}}+\mathcal{R}_{\textit{p}}+\mathcal{R}_{\textit{e}}$ 0.22 0.10 0.11 0.12 0.22 0.57 0.75
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{L}_{\textit{photo,local}}+\mathcal{R}_{\textit{p}}+\mathcal{R}_{\textit{e}}$ 0.21 0.10 0.11 0.10 0.23 0.59 0.77
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{L}_{\textit{photo,local}}+{\mathcal{R}_{\textit{s}}}+\mathcal{R}_{\textit{p}}+\mathcal{R}_{\textit{e}}$ 0.21 0.10 0.11 0.10 0.23 0.60 0.77
Table 3: Ablation study for our training losses.
As presented in Tab. 3, using only $\mathcal{L}_{\textit{geo}}$ yields suboptimal results. Combining $\mathcal{L}_{\textit{geo}}$ with $\mathcal{L}_{\textit{photo}}$ without regularizers performs poorly, indicating that integrating photometric consistency as objective term is non-trivial and can become unstable. The combination of $\mathcal{L}_{\textit{geo}}$ with $\mathcal{R}_{\textit{p}}$ significantly outperforms a simple ${\mathcal{R}_{\textit{s}}}$, demonstrating the importance of more sophisticated regularization. Adding the second mono depth map-based regularizer $\mathcal{R}_{\textit{e}}$, which aims to preserve or recover edges, further improves the results. Incorporating photometric consistency with regularizers notably boosts accuracy metrics, suggesting that local refinement with color loss is highly effective, though it does not improve coarse alignment. This is confirmed by the ablation studies where the photometric loss was applied only during local ($\mathcal{L}_{\textit{photo,local}}$), but not coarse, refinement. Based on these findings, we chose to use all regularizers and use the $\mathcal{L}_{\textit{photo}}$ only during local refinement.
Coarse & Local Refinement
Setup RMSE MAE L1-rel L1-inv $\delta_{1}$ Acc. Acc. Acc.
0.01 m 0.05 m 0.10 m
Coarse: none + Local: default 0.23 0.14 0.15 0.25 0.82 0.14 0.39 0.58
Coarse: $(o^{g},s^{g})$ + Local: default 0.16 0.08 0.08 0.13 0.91 0.24 0.61 0.77
[…]
Coarse: ${}_{\text{M}}$ + Local: default 0.20 0.09 0.10 0.14 0.92 0.25 0.64 0.80
[…]
Table 4: Ablation study for coarse and local refinement.
@@ 5.2 Ablation Studies
6 Limitations
While our approach demonstrates robust performance, several limitations remain.
Photometric consistency, though generally effective, can become a pitfall in scenarios involving specular materials, changing lighting conditions, or variations in white balance, leading to degraded results that our method will not mitigate.
As shown in Sec. 5.1, outliers in the sparse point cloud can result in spurious outcomes, affecting the overall accuracy and reliability of the depth estimation.
Finally, our method is dependent on the quality of the initial monocular depth map to establish a valid scene topology prior. This especially pertains to the availability of a good edge image to accurately mask out occluded regions which our losses rely on. Any deficiencies in these starting points can propagate through the subsequent steps, impacting the final results.
7 Conclusion
In conclusion, our approach demonstrates that monocular depth maps can be significantly improved by incorporating photometric consistency, leading to enhanced depth estimation. By utilizing a meshed version of the monocular depth map, we directly leverage the strong initial estimates and achieve stable optimization when handled correctly. In this context, our Poisson and edge regularizers have proven to be particularly effective, ensuring robust and reliable depth map refinement. Our simple yet effective pipeline consistently produces accurate and very dense depth maps, which is highly desirable for many downstream tasks.
8 Acknowledgments
Appendix A Implementation and Training Details
For our implementation, we use the hardware accelerated framework NVDiffRast [26] with the adjusted formulation outlined in Main-Sec. 4. For initial depth estimations we use Marigold [21] and employ COLMAP [40, 41] for Structure-from-Motion and image undistortion with configurations as recommended by the used dataset [67]. In Fig. 7, we depict sparse pointclouds of representative views, showcasing sparseness especially on white walls. For coarse refinement, $\Psi$ is a shallow MLP with degrees of positional encoding being $m=3$ and $k=5$. Two hidden layers of width 16, initialized with a normal distribution with $\sigma=0.1$. This comparatively small initialization ensures $o_{g}$ and $s_{g}$ start of very small, stabilizing training. During the coarse refinement phase, we used a learning rate of $0.001$ and stop after 400 iterations with batch sizes of 22. The current states of $o_{g}$ and $s_{g}$ are baked into $\mathbf{M}$’s $z$. We stop optimization after another 700 iterations of local refinement with learning rates of $0.0005$. In total, the process takes around 12 min per depth map but we see huge potential for performance improvements. The used losses were weighted as documented in Tab. 5. We used the Adam [24] optimizer with default parameters as proposed by the PyTorch framework [45].
Loss Regularizer
$\mathcal{L}_{\textit{photo}}$ 1.0 $\mathcal{R}_{\textit{p}}$ 400.0
$\mathcal{L}_{\textit{geo}}$ 0.1 $\mathcal{R}_{\textit{e}}$ 0.1
${\mathcal{R}_{\textit{s}}}$ 0.001
Table 5: Used weights for computing the total optimization loss.
Projection Matrices.
We convert the intrinsics definition $C$ with $\{f_{x},f_{y},c_{x},c_{y}\}$ to the projection matrix $P$ following the OpenGL convention [2], with $d_{\textit{near}}=0.1$ and $d_{\textit{far}}=100$ using
$P=\small\begin{pmatrix}\frac{2f_{x}}{W}&0&1-\frac{2c_{x}}{W}&0\\ 0&-\frac{2f_{y}}{H}&1-\frac{2c_{y}}{H}&0\\ 0&0&\frac{-(d_{\textit{far}}+d_{\textit{near}})}{d_{\textit{near}}-d_{\textit{far}}}&-\frac{2d_{\textit{far}}d_{\textit{near}}}{d_{\textit{far}}-d_{\textit{near}}}\\ 0&0&-1&0\end{pmatrix}.$ (19)
Appendix B Datasets
For the results present in Main-Sec. 5.1, we used 10 scenes from the ScanNet++ validation set and 10 depth maps of each (100 in total). These scenes have consecutive frames and less than 15 % of the views being marked as blurry. To tune the hyperparameters of our approach (see Main-Sec. 5.2), such as MLP configuration or decimation, we prepared another, different seven scenes with 10 views each (70 in total) in an analogous manner. Note that these seven scenes have not been used in the further evaluation. The hyperparameters found in the ablation studies were used consistently for all results showcased in the Main-Sec. 5.1.
For the Replica dataset, color and ground truth depth maps are obtained from the rendered meshes. We selected four scenes, and optimized 64 depth maps for each. As sparse point clouds, we generated 5000 point samples from the ground truth depth map, added a relative error with standard deviation of 2%, and randomly replaced 2% by randomly distributed outliers from the overall depth range. To examine the impact of the quality of the point cloud, we also generated worse point clouds: for two scenes (room0 and room2) we went down to 3500 points with relative error of 5% and 5% outliers, and for scenes office1 and office4 only 1000 points were generated, with a relative error of 10% and 10% outliers. The numbers were chosen to match approximately match the error distribution as observed in real-world scenarios, see Fig. 7. Nevertheless, the uniform density of the constructed sparse point clouds is unrealistic which may limit expressiveness.
To optimize for photo-consistency, we used 60 auxiliary views for the Replica dataset and 30 for Scannet++. MVSFormer, MVSFormer++ and GeoMVSNet received their default amount of 10 views as input which were evenly distributed across the set of our auxiliary views.
Appendix C Further Ablation studies

### v2

@@ Appendix C Implementation with GPU Acceleration
We convert the intrinsics definition $C$ with $\{f_{x},f_{y},c_{x},c_{y}\}$ to the projection matrix $P$ following the OpenGL convention [3], with $z_{\textit{near}}=0.1$ and $z_{\textit{far}}=100$ using homogeneous 4D coordinates for all vertices by
$P=\small\begin{pmatrix}\frac{2f_{x}}{W}&0&1-\frac{2c_{x}}{W}&0\\ 0&-\frac{2f_{y}}{H}&1-\frac{2c_{y}}{H}&0\\ 0&0&\frac{-(z_{\textit{far}}+z_{\textit{near}})}{z_{\textit{near}}-z_{\textit{far}}}&-\frac{2z_{\textit{far}}z_{\textit{near}}}{z_{\textit{far}}-z_{\textit{near}}}\\ 0&0&-1&0\end{pmatrix}$. (12)
This $P_{\{0..N\}}$ is then used in place of $C_{\{0..N\}}$ in Eq. 3.
@@
$\small z^{\prime}=\frac{z_{\textit{far}}+z_{\textit{near}}}{z_{\textit{near}}-z_{\textit{far}}}+\frac{2z_{\textit{far}}z_{\textit{near}}}{-z^{g}(z_{\textit{near}}-z_{\textit{far}})}.$ (13)
$u$ and $v$ are stored in normalized image coordinates in the [1,1][-1,1] range with
@@
An upside of this normalization is that it improves numerical stability and also simplifies positional MLP encoding (see Sec. 3.5).
@@
For both global and local refinement, depth optimization in reciprocal space leads to instable optimization. As such we do this optimization in linear space and substitute the rightmost part in Eq. 3 first by
@@
$\small\begin{array}[]{c}p^{\prime}=\min(-\frac{z_{\textit{far}}-\epsilon}{z},1)\max(-\frac{z_{\textit{near}}+\epsilon}{z},1)p^{\prime\prime}.\end{array}$ (16)
@@
Thus, Eq. 3 can be rewritten as:
$\small p^{*}_{i}=P_{i}[R_{i}|t_{i}][R_{0}|t_{0}]^{-1}p^{\prime},$ (17)
with $[R|t]$ matrices now being $4\times 4$ dimensional, respectively.
Tonemapping
Camera color ranges often vary with changing lighting, affecting the reliability of photometric losses. To simulate these variations, we use a method akin to analysis-by-synthesis by modeling the camera’s response curve and exposure settings. Consequently, we incorporated a differential tonemapping module, inspired by Rückert et al. [46] and common to NVS [18, 12].
In contrast to their implementation, our tonemapping function needs to be inverted. Thus, we opted for a trivially invertible gamma-curve instead of using a piecewise linear function to model the camera response curve.
Appendix D Dataset preparation
ScanNet++.
For the results present in Main-Sec. 4.2, we used 10 scenes from the ScanNet++ validation set and 10 depth maps of each (100 in total). These scenes were chosen to have 300 consecutive frames without cuts, as we chose auxiliary views based on the temporal neighborhood. Additionally, we only considered scenes with less than 15 % of the views being marked as blurry. To tune the hyperparameters of our approach (see Sec. F), such as MLP configuration or decimation, we prepared another, different seven scenes with 10 views each (70 in total) in an analogous manner. Note that these seven scenes have not been used in the evaluations of the main paper. The hyperparameters found in the ablation studies were used consistently for all results showcased in the Main-Sec. 4.2.
Replica.
For the Replica dataset, color and ground truth depth maps are obtained from the rendered meshes. We selected four scenes, and optimized over 64 views for each from the subset prepared by Rosinol et al. [45]. As sparse point clouds, we generated 5000 point samples from the ground truth depth map, added a relative error with standard deviation of 2%, and randomly replaced 2% by randomly distributed outliers from the overall depth range. To examine the impact of the quality of the point cloud, we also generated worse point clouds: for two scenes (room0 and room2) we went down to 3500 points with relative error of 5% and 5% outliers, and for scenes office1 and office4 only 1000 points were generated, with a relative error of 10% and 10% outliers. The numbers were chosen to approximately match the error distribution as observed in real-world scenarios, see Fig. 8. Nevertheless, the uniform density of the constructed sparse point clouds is unrealistic which may limit expressiveness.
Number of views.
To optimize for photo-consistency, we used 60 auxiliary views for the Replica dataset and 30 for ScanNet++. Auxiliary views were simply the set of consecutive frames which are temporally closest to $\mathbf{I}_{0}$. MVSFormer, MVSFormer++ and GeoMVSNet received their default amount of 10 views as input which were evenly distributed across the set of our auxiliary views.
Note that our approach works best with a high number of auxiliary views with a distribution that can be considered as a front facing capture, while input views with extreme baselines offer little benefits due to the masking of regions with potential disocclusions. A higher number of auxiliary views will generally increase robustness and accuracy of the optimization. The number of used input views was thus empirically set according to the estimated framerate of the provided dataset, covering approximately a 1-2 s video clip. (Replica has approximately 30 frames per second, ScanNet++  10 - 15, exact timestamps are, to the best of our knowledge, not provided).
Tanks and Temples.
Regarding the Tanks and Temples results, we used every 5th frame of the original video for Ignatius, every 15th frame for Courthouse and every 30th frame for Truck. The views for our multi-view optimization were simply selected by using the 30, 15, and 7 preceding and succeeding frames (for Ignatius, Courthouse, and Truck respectively). We used the same view selection for MVSFormers. For point cloud fusion, we stuck to the implementation by Cao et al. [4] (MVSFormer). This implementation receives a confidence map and a respective threshold as input, which we set to 0.8. To fuse our pointclouds, we used the edge map $\mathbf{E}$ as confidence map.
Appendix E Additional Results
Qualitative results on Replica.
Visual examples of Replica results are shown in Fig. 9. Noise like artifacts (see e.g the pillow case and wall in the center column) can emerge when there are severe outliers in the sparse point cloud as in the case of ’Ours (-)’. Generally, our method produces dense and clean results also for feature scarce areas, where others struggle.
Figure 9: Results on the Replica dataset with MVSFormers’ certainty threshold at their default of 0.5. Green borders indicate image crops, while red highlights error maps; brighter areas imply greater errors, and grey indicates missing samples.
Full scene reconstruction on Tanks and Temples.
Method Ignatius  (t= 0.003) Courthouse  (t= 0.025) Truck  (t= 0.005)
   Metric $\uparrow$ (P)recision (R)ecall (F)-Score P R F P R F
MVSFormer 0.59 0.97 0.74 0.48 0.60 0.53 0.74 0.59 0.66
MVSFormer++ 0.62 0.96 0.75 0.51 0.59 0.55 0.68 0.78 0.73
COLMAP 0.69 0.72 0.71 0.25 0.47 0.33 0.62 0.64 0.63
Ours 0.55 0.86 0.67 0.38 0.48 0.43 0.50 0.54 0.52
@@ Appendix E Additional Results
Additionally, to individual depth map comparisons, we evaluated on the task of full scene reconstruction which implies a fusion and filtering process of involved depth maps. In Tab. 5, we present the scores of the dense point cloud reconstructions following the standard evaluation protocol of the Tanks and Temples reconstruction benchmark [28] and visual results are presented in Fig. 10.
We used COLMAP’s default settings for the dense reconstruction used the resulting SfM poses for all other methods. For generating both MVSFormers’ and our results, we employed Yao et al.’s point cloud fusion method [71].
Again we used the same configuration and hyperparameters as in the main paper to produce our results, with two exceptions: We changed the learning rate (1e-3 $\rightarrow$ 5e-4) during coarse alignment of the Truck scene. In this case, the scene scale was very different compared to the other scenes. Secondly, we applied the tonemapper presented in the main paper to account for challenging exposure changes. An ablation study is presented in Suppl.-Sec. F. We incorporate results using both Marigold and DepthAnything2 as input to our method, since we found that Marigold usually performs better indoors, while DepthAnything2 is favorable for outdoor settings.
We observe that our method performs worse on the Tanks and Temples dataset compared to the state-of-the-art. In contrast to MVSFormer and MVSFormer++, which are mainly trained on semantically related scenes, our method was not adapted for or tuned to the Tanks and Temples dataset and outperforms the same methods on the other datasets. In addition, Tanks and Temples only provides point clouds as ground truth and therefore many advantages of our method, like high completeness per depth map, are not beneficial in this evaluation.
Figure 10: Qualitative results of the Tanks and Temples Scene Ignatius. Brighter is better
Appendix F Ablation Studies
In this section, we evaluate different choices of hyper parameter settings and present ablation studies based on seven withheld scenes of the ScanNet++ dataset, referred to as the ablation or withheld scene set.
Loss Ablations.
[…]
$\downarrow$ 0.01 m 0.05 m 0.10 m
$\scriptstyle\mathcal{L}_{\textit{geo}}$ 0.32 0.15 0.27 0.13 0.14 0.47 0.69
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{L}_{\textit{photo}}$ 0.31 0.15 0.14 0.14 0.16 0.42 0.59
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{R}_{\textit{p}}$ 0.24 0.11 0.12 0.11 0.18 0.55 0.76
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{R}_{\textit{p}}+\mathcal{R}_{\textit{e}}$ 0.21 0.10 0.11 0.10 0.17 0.57 0.76
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{L}_{\textit{photo,(c\&l)}}+\mathcal{R}_{\textit{p}}+\mathcal{R}_{\textit{e}}$ 0.18 0.08 0.08 0.09 0.27 0.64 0.80
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{L}_{\textit{photo}}+\mathcal{R}_{\textit{p}}+\mathcal{R}_{\textit{e}}$ 0.19 0.09 0.08 0.09 0.27 0.64 0.80
Table 6: Ablation study on used training losses, with $\mathcal{L}_{\textit{photo,(c\&l)}}$ indicates applying photometric loss also in coarse refinement instead of only local refinement. Our default configuration is in blue.
As presented in Tab. 6, using only $\mathcal{L}_{\textit{geo}}$ (see Suppl. Sec. A) performs subpar. When $\mathcal{L}_{\textit{geo}}$ gets combined with the proposed regularizers, distance metrics like RMSE, MAE, or L1-rel are improved. Adding $\mathcal{L}_{\textit{photo}}$ during local refinement, notably boosts accuracy measures (30 % relative improvement in case of $\tau$=0.01 m and 11 % in case of 0.05 m). Furthermore, as seen in Fig. 11, photometric loss is able to finely adjust inaccuracies in the depth map.
While aiming for photometric consistency also during coarse refinement (indicated by $\mathcal{L}_{\textit{photo,(c\&l)}}$) increases the risk of getting stuck in local minima, this proved irrelevant in the case of the ablation scene set.
Figure 11: Results of depth maps under warping. ’w/ photo’ indicates use of $\mathcal{L}_{\textit{photo}}$, ’w/o photo’ omits this. Error maps shown on the right; brighter areas imply greater errors.
Photometric Loss Choice.
Ablation on the withheld scene set, extends Tab. 6. while invariances of the VGG loss did not guide towards precise local matching, adding a perceptual loss like SSIM performed on par or improved metrics, underlining the effectiveness of photometric consistency as an objective function.
[…]
$\downarrow$ 0.01 m 0.05 m 0.10 m
Default losses 0.18 0.08 0.08 0.09 0.27 0.64 0.80
+SSIM 0.18 0.08 0.10 0.10 0.36 0.69 0.82
+VGG 0.25 0.13 0.14 0.15 0.18 0.48 0.65
Table 7: Configuration used to produce results of the main paper is marked in blue.
Alternative Regularizers.
Configuration RMSE MAE L1-rel L1-inv $\uparrow Acc.$ Acc.
$\downarrow$ 0.01 m 0.05 m 0.10 m
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{L}_{\textit{photo}}+{\mathcal{R}_{\textit{s}}}$ 0.391 0.194 0.296 0.161 0.152 0.414 0.582
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{L}_{\textit{photo}}+\mathcal{R}_{\textit{n}}$ 0.188 0.080 0.080 0.088 0.193 0.589 0.805
$\scriptstyle\mathcal{L}_{\textit{geo}}+\mathcal{L}_{\textit{photo}}+\mathcal{R}_{\textit{p}}+\mathcal{R}_{\textit{e}}$ 0.177 0.082 0.077 0.088 0.273 0.644 0.801
Table 8: Ablation study on used smoothing or gradient-based regularizers. Ours is marked in blue.
Tab. 8 shows results for replacing our Poisson regularizer $\mathcal{R}_{\textit{p}}$ and edge-aware regularizer $\mathcal{R}_{\textit{e}}$ by a very simple smoothness term $\mathcal{R}_{\textit{s}}$} or using an estimated normal map [1] for regularization indicated by $\mathcal{R}_{\textit{n}}$.
The $\mathcal{R}_{\textit{s}}$} is a simple regularizer that suppresses noisy results:
[…]
where $g$ is a Gaussian blur with window size 5.
$\mathcal{R}_{\textit{n}}$ measures the cosine similarity of the Sobel-based normal of $\mathbf{D}^{*}_{0}$ with the estimate of the normal predictor [1].
Our choice of regularizers significantly outperforms $\mathcal{R}_{\textit{s}}$} demonstrating the importance of more sophisticated regularization. Simultaneously, our regularizers are on par with $\mathcal{R}_{\textit{n}}$.
Coarse & Local Refinement
Setup RMSE MAE L1-rel $\delta_{1}$ Acc. Acc. Acc.
$\downarrow$ $\uparrow$ 0.01 m 0.05 m 0.10 m
Coarse: none; Local: default 0.23 0.14 0.15 0.82 0.14 0.39 0.58
Coarse: $(o^{g},s^{g})$; Local: default 0.16 0.08 0.08 0.91 0.24 0.61 0.77
[…]
Table 9: Ablation study for coarse and local refinement. Our used configuration is highlighted is blue.
[…]